If a perimeter is 130 feet and the length is 5 feet longer than the width, then what is the length and width?
Answer by macston(5194)

.
W=width; L=length=W+5 feet; P=perimeter=130 feet
P=2(L+W) Substitute for L
130 feet=2(W+5 feet+W) Divide each side by 2.
65 feet=2W+5 feet Subtract 5 feet from each side.
60 feet=2W Divide each side by 2.
30 feet=W ANSWER 1: The width is 30 feet
L=W+5 feet=30 feet+5 feet=35 feet ANSWER 2: The length is 35 feet.
.
CHECK:
.
P=2(L+W)
130 feet=2(35 feet+30 feet)
130 feet=2(65 feet)
130 feet=130 feet